OpenStudy (anonymous):

5 to the -2 power = ? Please help! I am a 7th grader so if anyone could not just give me the answer but explain how you got the answer mwould be great. Thank you all!

OpenStudy (anonymous):

Well, here we go: so anything (x) to the negative 1 power is the same as 1/x: Example: \[4^{-1}\] = 1/4 and \[23^{-1}\] = 1/23: But what happens is, if it;s to the - n power, it becomes 1/x^n: Ex: \[4^{-2}\] = 1/4^2 = 1/16: So, using your example: \[5^{-2} = \frac{ 1 }{ 5^2 } = \frac{ 1 }{ 25 }\]

OpenStudy (anonymous):

Thank you so much. Have a great day